OpenStudy (raffle_snaffle):

In January in Montana, the temperature on the Celsius scale reads identical to the temperature on the Fahrenheit scale. (a) What is this temperature? (b) What is this temperature in K? (c) What is this temperature in °R?

OpenStudy (raffle_snaffle):

This question is confusing? Does she literally want me to look up the temp in Montana in Jan and use it or solve using the formulas?

OpenStudy (radar):

No, there is a temperature that C and F are equal. That temperature is -40 degrees. Look here: https://www.quora.com/What-is-the-reason-40-Fahrenheit-the-same-as-40-Celsius-Hows-the-conversion-work-to-get-the-same-number-at-40

OpenStudy (radar):

See how the graphs will intersect at -40, the two equations are linear.

OpenStudy (osprey):

In kelvins its 273 plus celsius. So, freezing in celsius is 273 K. room temp is 20+273 =293. unobtainable absolute zero is -273 celsius, and 0 Kelvins. There seem to be several scales here Kelvins, Rankines, Fahrenheits, and Celsius. I think that Rankines refers to the Rankine cycle in heat engines thermodynamics, Kelvins is to do with Cryogenics, Celsius is to do with us and the weather, and Herr Fahrenheit is the "old" version of "reasonable" temperature measurements. There are formulae which link these "metrics" are they called ? and maybe there's a factor of 9/5 in F to C or vice versa. But, basically, if you are hot at, say, 30 celsius, then you may like to know that you could be over 100 fahr... And I think that medical body temperatures may still be quoted in Fs. So, a temperature of 104 F is a fever type territory ? Best wishes to Montana - cold there is it ?

OpenStudy (osprey):

I also think that most temperature scales can be calibrated against a thermodynamic temperature scale, which is to do with heat engines and hot and cold reservoirs, and possibly helps to explain the mayhem !!! Bon voyage a vous

OpenStudy (osprey):

Robert Frost, Engineer with specialization in spacecraft operations, orbital mechanics, and... I've just looked up the link given in one of the posts. I love that surname !

OpenStudy (radar):

Take one of the conversion equations: Tf = 9/5 Tc + 32, then assume the temperatures are equal at temperature T, substituting T for Tf and Tc you now have: t = (9/5)t + 32 Solving for t: t - (9/5)t = 32 -(4/5)t = 32 t = -40

OpenStudy (raffle_snaffle):

radar that is what I kind of initially did.

OpenStudy (radar):

Great, I guess the point of the problem, stay out of Montana in January.